Q. Find the wrong number in the series
8, 17, 37, 79, 165, 338

Answer by fcabanski(1391)

Each number is progressively 1, 3, 5, 7... more than twice the previous number.


8+8 = 16 + 1 = 17


17+17 = 34 + 3 = 37


37 + 37 = 74 + 5 = 79


79 + 79 = 158 + 7 = 165


165 + 165 = 330+ 9 = 339. The final number should be 339.